# The Devil behind the mask: An emergent safety vulnerability of Diffusion LLMs

### v1

@@ 4.2 Main Results
Figure 4: Illustrative cases of harmful completions generated by four dLLMs when attacked by DiJA. The red text represents harmful contents generated by dLLMs under DiJA attack.
Despite exhibiting safety on par with autoregressive models, dLLMs remain highly vulnerable to our proposed automatic diffusion-based LLM jailbreak attack pipeline, DiJA. Experimental results of our jailbreak attacks are presented in Tables 12, and 3. Specifically, our proposed DiJAachieves surprisingly strong attack performance across three jailbreak benchmarks, an outcome that is rarely observed on autoregressive models. This is because our method exposes the harmful intent in the prompt directly, without any rewriting, obfuscation, or decomposition. (i) For keyword-based ASR (ASR-k), we consistently achieved the highest attack success rates across all benchmarks on all evaluated dLLMs, with some models even reaching a 100% success rate. (ii) On Dream-Instruct, the safest dLLM among the four evaluated, our evaluator-based ASR (ASR-e) on HarmBench surpasses that of the second-best method, ReNeLLM, by 54%. On JailbreakBench, the improvement reaches 78.5%, and on StrongREJECT, our SRS exceeds ReNeLLM’s by 37.7. (iii) We observe that using GPT-4o (i.e., DiJA*) yields a slight advantage in attack effectiveness compared to using Qwen-2.5-7B-Instruct (i.e., DiJA). Upon inspection, we attribute this to GPT-4o’s superior few-shot in-context learning and instruction-following capabilities.
@@
Furthermore, to assess the robustness of DiJA, we evaluate its attack performance under two commonly used defense methods, Self-reminder (Xie et al., 2023) and Robust Prompt Optimization (Zhou et al., 2024). For detailed information regarding the defenses, please refer to the Appendix A.4. As demonstrated in Figure 5, despite the application of defense methods, our proposed attack maintains its effectiveness with no significant degradation, highlighting the robustness of our approach.
@@ 5 Analysis
The Figure 6 illustrates the differential behavior of the LLaDA-1.5 among dLLMs in response to the same sensitive query under varying generation lengths. At shorter generation lengths, the dLLM consistently refuses to engage with the query, adhering to ethical guidelines by responding with “Sorry, but I can’t assist with that.” However, at longer generation lengths, the dLLM initially acknowledges the sensitive nature of the topic but proceeds to provide content that could be considered harmful. We hypothesize that the dLLM’s bidirectional attention mechanism, which allows it to consider future tokens during token generation, influences its response based on the allowed generation length. Shorter lengths seem to prompt more conservative, ethically aligned responses, while longer lengths could result in more detailed, albeit potentially harmful, content generation.
@@
Inspired by this, we also investigate the effect of the number of masked tokens—analogous to generation length—in our proposed DiJA framework. Specifically, we adopt a regular-expression-based approach to replace every masked token segment in the context-aware interleaved mask-text jailbreak prompts with a specified number of masks. We then evaluate the impact of varying the number of masked tokens on attack performance across JailbreakBench and StrongREJECT. As shown in Figure 7, when the number of masked tokens is relatively small (e.g., 10), the attack effectiveness across all four dLLMs is limited. This is consistent with the observation in Figure 6, where a small generation length makes it difficult to elicit harmful content. Meanwhile, we also observe that when the number of masked tokens becomes too large (e.g., 50), the attack success rate, StrongREJECT score, and harmfulness score tend to decline in many cases. We find that decoding an excessive number of masked tokens can lead to overly long generations, which often contain irrelevant or meaningless content. This, in turn, may negatively impact the attack effectiveness scores.
@@ 6 Conclusion
Appendix A More Implementation Details
A.1 Victim Models

### v2

@@ 4.2 Main Results
Figure 4: Illustrative cases of harmful completions generated by four dLLMs when attacked by DiJA. The red text represents harmful content generated by dLLMs under DiJA attack.
Attack Effectiveness. Despite exhibiting safety on par with autoregressive models, dLLMs remain highly vulnerable to our proposed automatic diffusion-based LLM jailbreak attack pipeline, DiJA. Experimental results of our jailbreak attacks are presented in Tables 12, and 3. Specifically, our proposed DiJA achieves surprisingly strong attack performance across three jailbreak benchmarks. This is because our method exposes the harmful intent in the prompt directly [^9] Here, “direct exposure” refers to cases where the original harmful instruction is fully preserved, as illustrated in Figure 4., without any rewriting, obfuscation, or decomposition, nor requiring role-playing, deceptive scenario nesting, or other indirection. (i) For keyword-based ASR (ASR-k), we consistently achieved the highest attack success rates across all benchmarks on all evaluated dLLMs, with some models even reaching a 100% success rate. (ii) On Dream-Instruct, the safest dLLM among the four evaluated, our evaluator-based ASR (ASR-e) on HarmBench surpasses that of the second-best method, ReNeLLM, by 54%. On JailbreakBench, the improvement reaches 78.5%, and on StrongREJECT, our SRS exceeds ReNeLLM’s by 37.7. (iii) We observe that using GPT-4o (i.e., DiJA*) yields a slight advantage in attack effectiveness compared to using Qwen-2.5-7B-Instruct (i.e., DiJA). Upon inspection, we attribute this to GPT-4o’s superior few-shot in-context learning and instruction-following capabilities.
@@
Robustness Against Defenses. Furthermore, to assess the robustness of DiJA, we evaluate its attack performance under two commonly used defense methods, Self-reminder (Xie et al., 2023) and Robust Prompt Optimization (Zhou et al., 2024). For detailed information regarding the defenses, please refer to the Appendix C.4. As demonstrated in Figure 5, despite the application of defense methods, our proposed attack maintains its effectiveness with no significant degradation, highlighting the robustness of our approach.
Preliminary Alignment for dLLMs. Spurred by these vulnerabilities, we pilot an architecture-aware defense: Refusal-Aware Denoising Alignment, which teaches dLLMs to spot interleaved mask-text jailbreak patterns and firmly decline unsafe requests. As shown in Table 4 (details in Appendix A), even this early attempt significantly boosts rejection of malicious prompts, hinting that robust safety alignment for dLLMs is not only possible but calls for tailored, architecture-specific strategies.
@@ 5 Analysis
The Figure 6 illustrates the differential behavior of the LLaDA-1.5 among dLLMs in response to the same sensitive query under varying generation lengths. At shorter generation lengths, the dLLM consistently refuses to engage with the query, adhering to ethical guidelines by responding with “Sorry, but I can’t assist with that.” However, at longer generation lengths, the dLLM initially acknowledges the sensitive nature of the topic but proceeds to provide content that could be considered harmful. We hypothesize that the model’s bidirectional attention mechanism, which allows it to consider future tokens during token generation, influences its response based on the allowed generation length. Shorter lengths tend to elicit more conservative, ethically aligned responses, whereas longer lengths can yield more detailed and potentially harmful content.
@@
Inspired by this observation, we investigate the effect of number of masked tokens, which is analogous to generation length, in our proposed DiJA. Specifically, we adopt a regular-expression-based approach to replace every masked token segment in the context-aware interleaved mask-text jailbreak prompts with a specified number of masks. We then evaluate the impact of varying the number of masked tokens on attack performance across JailbreakBench and StrongREJECT. As shown in Figure 7, when the number of masked tokens is relatively small (e.g., 10), the attack effectiveness across all four dLLMs is limited, consistent with the observation in Figure 6 that short generation lengths make it difficult to elicit harmful content. Meanwhile, when the number of masked tokens becomes too large (e.g., 50), the attack success rate, StrongREJECT score, and harmfulness score often decline. Decoding an excessive number of masked tokens can lead to overly long generations that contain irrelevant or meaningless content, which in turn may negatively impact attack effectiveness.
@@ 6 Conclusion
Ethics statement
Our research identifies a significant safety vulnerability in diffusion-based large language models (dLLMs) and, in response, proposes and validates targeted defense and alignment solutions. Our goal is to proactively improve AI safety. We recognize the dual-use nature of our work but believe the benefits of disclosing this vulnerability to catalyze countermeasures outweigh the risks of potential misuse. Our research was conducted with integrity in a controlled environment to foster safer AI development, and we do not condone the use of our methods to cause harm.
Reproducibility statement
We aim to make our work fully reproducible. The core algorithmic ideas and assumptions are detailed in Section 3, with a step-by-step pseudocode in Algorithm 1. Experimental settings, including model lists and benchmark coverage, are summarized in Section 4 and expanded in Appendix C (victim models, benchmarks, attack baselines, and hyperparameters). Evaluation protocols and exact judge prompts are specified in Appendix D (including Tables 19 and 20); defense configurations and prompts are in Appendix C.4 (Tables 13 and 14). Our in-context learning template and illustrative prompt constructions for DiJA are provided in Appendix C.5 (Table 15). Additional results, ablations, and analyses appear in Appendix B. We will release our codebase, including evaluation scripts, data processing utilities, alignment data, aligned model weights, and training recipes, to facilitate the reproduction of our results and findings.
Acknowledgements
This paper was sponsored by CCF-Tencent Open Research Fund.
@@ Acknowledgements
Appendix A Defense and Alignment for dLLMs: A Preliminary Exploration
Having demonstrated the emergent safety vulnerabilities in dLLMs stemming from their core architectural properties, a critical next step is to investigate whether these models can be effectively aligned to resist such attacks. Standard safety protocols, primarily designed for autoregressive LLMs, are ill-suited for dLLMs because they fail to account for inherent architectural properties like bidirectional context modeling and parallel decoding, which our DiJA framework exploits. In this section, we present a preliminary exploration into a novel, architecture-specific defense strategy designed to mitigate these newly identified risks.
A.1 Strategy: Refusal-Aware Denoising Alignment
Our core hypothesis is that instead of preventing the model from processing a malicious prompt (i.e., an interleaved mask-text prompt), we can train it to recognize the adversarial structure of an interleaved mask-text prompt and respond with a firm but safe refusal. We term this approach Refusal-Aware Denoising Alignment. The objective is to fine-tune the dLLM to associate the specific pattern of an interleaved mask-text jailbreak prompt not with a contextually coherent harmful completion, but with a pre-defined safety response. This effectively teaches the model a new, safe behavior for the denoising process when confronted with a known attack vector.
This strategy is conceptually aligned with recent advancements in “Deep Alignment” for autoregressive models, such as the work by (Qi et al., 2024). They observed that standard alignment is often “shallow” and vulnerable to prefilling attacks that force the model into a harmful state. To counter this, they proposed training interventions that teach the model to revert to a refusal even after a harmful prefix. Similarly, our Refusal-Aware Denoising Alignment can be viewed as an adaptation of this deep alignment principle to the dLLM architecture. Since dLLMs do not generate sequentially from a prefix but rather denoise globally based on bidirectional context, standard refusal training fails when the context (via interleaved masks) implies compliance. Our method extends the deep alignment philosophy by training the dLLM to recognize these adversarial structural patterns and enforcing a refusal trajectory, effectively deepening the alignment to persist even when the prompt context attempts to bypass initial safety guardrails.
A.2 Methodology: Curating a Targeted Alignment Dataset
To implement this strategy, we constructed a specialized alignment dataset. The process is as follows:
  1. 1.
    Data Sourcing: We began with a corpus of approximately 10,000 harmful instructions, combining 5,000 prompts from the WildGuard dataset (Han et al., 2024) with 5,000 prompts from Circuit Breaker Dataset (Zou et al., 2024).
  2. 2.
    Adversarial Prompt Generation: We processed these harmful prompts through our DiJA pipeline to generate their corresponding interleaved mask-text jailbreak prompts. This produced the adversarial input that the model needs to learn to identify.
  3. 3.
    Refusal Pairing: In the crucial step, instead of generating harmful content for the masked sections, we systematically replaced the expected malicious output with a standardized refusal message, such as “I’m sorry, I can’t help with that.” This creates a direct pairing between the full, unaltered DiJA prompt and a safe refusal response. The final dataset is thus a collection of prompt-refusal alignment pairs, where each pair consists of an adversarial interleaved mask-text prompt and its corresponding desired safe refusal.
A.3 Implementation and Training
We fine-tune LLaDA-8B-Instruct (Nie et al., 2025b) on the curated alignment corpus to explicitly remap the denoising behavior of masked spans that follow adversarial interleaved prompts: from context-preserving completion to safety-aligned refusal. During training, the input is the full adversarial interleaved mask-text prompt that contains the malicious instruction and benign separators. A standardized refusal is appended to the prompt, and the refusal tokens are masked in the model state. The optimization objective is to reconstruct (denoise) the masked refusal across diffusion steps while keeping unmasked tokens fixed, so that the model learns to emit a refusal when this adversarial structure is present.
Table 4: Jailbreak results of DiJA on HarmBench, JailBreakBench, and StrongREJECT, comparing LLaDA-Instruct with LLaDA-Instruct-Aligned.
Benchmarks HarmBench JailBreakBench StrongREJECT
Metrics ASR-k ASR-e HS ASR-k ASR-e HS ASR-k SRS HS
LLaDA-Instruct 96.3 55.5 4.1 95.0 81.0 4.6 92.7 60.8 4.7
[…]
Table 5: Comparison between vanilla LLaDA-Instruct and LLaDA-Instruct-Aligned on benign general benchmarks (GSM8K, GPQA, BBH, HumanEval, MBPP, MMLU-Pro, MMLU).
Models GSM8K GPQA BBH HumanEval MBPP MMLU-Pro MMLU
LLaDA-Instruct 78.5 32.4 51.5 31.7 39.2 35.1 65.7
LLaDA-Instruct-Aligned 78.9 27.5 47.8 29.8 37.3 36.1 64.2
A.4 Results and Implications
As shown in Table 4, our preliminary results indicate that this targeted fine-tuning significantly enhances the model’s ability to reject harmful instructions delivered via the interleaved mask-text jailbreak prompts. This finding is a crucial first step, demonstrating that dLLMs are not inherently un-alignable. However, it underscores that their alignment requires bespoke, architecture-aware strategies. Simply applying safety alignment techniques from the autoregressive domain is insufficient. Additionally, we evaluated the aligned dLLM on general benchmarks, with results shown in Table 5. Despite fine-tuning on only 10,000 alignment examples, which can reasonably be expected to incur some performance degradation, since no prior training data were mixed in to preserve capabilities, the aligned model’s performance on general benchmarks remains well within an acceptable range. Taken together with Table 4, these findings show that we obtain a substantial improvement in safety at only a modest (and likely further reducible) cost to general capability. Looking ahead, we defer a detailed discussion of next steps on dLLMs alignment to Section E. We will also release the alignment dataset, training code, and aligned model weights to facilitate reproducibility.
Table 6: Results on code-oriented dLLMs under the HarmBench. We report ASR-k, ASR-e, and HS for DiffuCoder-7B-Instruct, DiffuCoder-7B-cpGRPO, and Dream-Coder-v0-Instruct. DiJA (Ours) shows a marked improvement over Zeroshot.
Victim Models DiffuCoder-7B-Instruct DiffuCoder-7B-cpGRPO Dream-Coder-v0-Instruct
Metrics ASR-k ASR-e HS ASR-k ASR-e HS ASR-k ASR-e HS
Zeroshot 67.8 22.0 2.9 53.0 14.5 2.2 75.0 30.8 3.5
[…]
Table 7: Results on code-oriented dLLMs under the JailbreakBench. We report ASR-k, ASR-e, and HS for DiffuCoder-7B-Instruct, DiffuCoder-7B-cpGRPO, and Dream-Coder-v0-Instruct. DiJA (Ours) shows a marked improvement over Zeroshot.
Victim Models DiffuCoder-7B-Instruct DiffuCoder-7B-cpGRPO Dream-Coder-v0-Instruct
Metrics ASR-k ASR-e HS ASR-k ASR-e HS ASR-k ASR-e HS
Zeroshot 41.0 29.0 2.6 34.0 12.0 1.7 32.0 37.0 2.8
[…]
Table 8: Results on code-oriented dLLMs under the StrongREJECT Benchmark. We report ASR-k, SRS, and HS for DiffuCoder-7B-Instruct, DiffuCoder-7B-cpGRPO, and Dream-Coder-v0-Instruct. DiJA (Ours) shows a marked improvement over Zeroshot.
Victim Models DiffuCoder-7B-Instruct DiffuCoder-7B-cpGRPO Dream-Coder-v0-Instruct
Metrics ASR-k SRS HS ASR-k SRS HS ASR-k SRS HS
Zeroshot 38.3 9.7 2.3 30.4 6.2 1.6 53.9 14.6 3.3
[…]
Appendix B More Experimental Results and Analysis
B.1 DiJA Attack on Code-Oriented dLLMs
As the promise of dLLMs for coding becomes increasingly clear and a wave of code-focused dLLMs continues to emerge (Labs et al., 2025; Gong et al., 2025; Xie et al., 2025), these models may see broad adoption due to their strengths in code generation. It is therefore essential to examine their vulnerabilities to attack, which also offers a further validation of DiJA’s effectiveness. To this end, we conducted experiments on three open-source code dLLMs: DiffuCoder-7B-Instruct, DiffuCoder-7B-cpGRPO, and Dream-Coder-v0-Instruct. As shown in Tables 6, 7, and 8, our method substantially compromises the defenses of these code dLLMs, leading to marked increases across attack metrics. This should command serious attention: it is imperative to apply architecture-aware alignment during training to avert significant safety risks in real-world deployments.
B.2 Comprehensive Evaluation of dLLMs Defense Capabilities
Furthermore, we comprehensively evaluated the defensive capabilities of dLLMs and autoregressive LLMs against various jailbreak attacks using three benchmarks: HarmBench, JailBreakBench, and StrongREJECT. As shown in Figure 8, dLLMs exhibit an evaluator-based ASR (ASR-e) that is comparable to or even lower than that of autoregressive LLMs under the AIM and PAIR attacks. This trend is consistent with the findings presented in Figure 3 in the main text. Meanwhile, the results in Figure 9 indicate that dLLMs generally exhibit comparable or slightly better initial resistance to zero-shot attacks compared to autoregressive LLMs, as evidenced by lower keyword-based and evaluator-based attack success rates (ASR-k and ASR-e). In summary, our comprehensive evaluation across multiple benchmarks and attack scenarios reveals that dLLMs often match or surpass those of autoregressive LLMs in resisting existing jailbreak attack methods.
B.3 Systematic Analysis of Prompt Diversity
We conducted a systematic quantitative analysis for the diversity of the interleaved mask-text prompts generated by DiJA. Since DiJA transforms vanilla harmful queries into interleaved mask-text formats while preserving the original malicious intent, diversity is primarily manifested in the masking patterns rather than semantic shifts. In our framework, the masking patterns mainly include three types: block-wise masking, fine-grained masking, and progressive masking.
We evaluate this structural diversity using two key metrics on a representative set of 400 interleaved mask-text prompts generated from the HarmBench dataset:
  • Mask Ratio: The percentage of masked tokens relative to the total token count per prompt. This metric reflects the overall density of information concealment.
  • Average Mask Span Length: The average length of contiguous masked segments within a prompt. This metric differentiates between fine-grained deletions (short spans) and block-wise redactions (long spans).
Table 9: Statistical analysis of structural diversity metrics for DiJA prompts on HarmBench ($N=400$). The wide range and variance indicate that our method successfully generates a diverse array of masking patterns, covering both fine-grained and block-wise structures.
Metric Min Max Mean Std. Dev
Mask Ratio (%) 15.3 78.6 62.1 12.4
Avg. Mask Span Length (tokens) 7.4 65.8 28.5 18.2
As shown in Table 9, our analysis reveals a broad and healthy distribution across both metrics. The Mask Ratio spans a wide range from 15.3% to 78.6%, indicating that our generated prompts vary from lightly edited contexts (requiring minor infilling) to heavily masked templates (requiring substantial generation). Similarly, the Average Mask Span Length exhibits significant variance, ranging from short spans of $\sim$7.4 tokens (characteristic of Fine-grained Masking) to long blocks of $\sim$65.8 tokens (characteristic of Block-wise Masking). This quantitative evidence confirms that DiJA does not converge on a single, repetitive template. Instead, by leveraging the strategies described in Section 3, it produces a diverse array of adversarial structures that effectively probe different failure modes of the dLLM’s bidirectional attention mechanism.
B.4 Reproducibility and Stability Analysis
Since DiJA utilizes an auxiliary LLM (e.g., Qwen2.5-7B-Instruct) to generate adversarial interleaved mask-text prompts via in-context learning, the specific structure of the generated prompts and the number of mask tokens may vary across different generation runs. To verify the reproducibility and stability of our method against this generation stochasticity, we conducted a systematic stability analysis.
Specifically, we performed three independent runs of the DiJA attack pipeline on the HarmBench dataset. To introduce sufficient variance, we applied different random seeds and varied the temperature settings ($T\in\{0.2,0.3,0.4\}$) for the auxiliary prompt generation model (Qwen2.5-7B-Instruct) across the runs.
Table 10: Reproducibility analysis of DiJA on HarmBench across three independent runs with varying temperature settings. The low standard deviation indicates high stability.
Metrics ASR-k ASR-e HS ASR-k ASR-e HS ASR-k ASR-e HS ASR-k ASR-e HS
Run 1 ($T=0.2$) 96.3 55.5 4.1 95.8 56.8 4.1 98.3 57.5 3.9 97.5 46.8 3.9
Run 2 ($T=0.3$) 95.8 55.0 4.0 95.3 56.3 4.0 97.5 56.8 3.8 96.8 46.3 3.9
Run 3 ($T=0.4$) 96.8 56.3 4.2 96.3 57.3 4.2 98.8 58.0 3.9 98.0 47.3 3.8
[…]
The quantitative results are summarized in Table 10. As demonstrated, DiJA maintains highly consistent performance across repeated independent runs, with the standard deviation of ASR remaining consistently low ($0.7\%$). This stability is rooted in our design philosophy: DiJA exploits the fundamental architectural characteristics of dLLMs (i.e., bidirectional context modeling and parallel decoding) rather than relying on brittle, stochastic heuristics or specific prompt artifacts. This robustness further confirms that the safety vulnerability we exposed is intrinsic to the current dLLM paradigm and underscores the urgent need for effective defense solutions.
B.5 Ablation Study
To systematically isolate the contributions of individual components within the DiJA framework and verify that the attack’s effectiveness stems from the proposed masking mechanism rather than auxiliary factors (e.g., LLM-based refinement), we conducted a comprehensive ablation study.
Specifically, we evaluated three variants on the HarmBench dataset using LLaDA-8B-Instruct as the victim model:
  1. 1.
    Ablation on Prompt Refinement (w/o Refinement LLM): To verify the high attack success rate is not merely due to refinement of the auxiliary LLM, we completely excluded the LLM and the diverse few-shot demonstrations. Instead, we used a fixed, heuristic template for benign separators (e.g., “First, [MASK]… Second, [MASK]… Third, [MASK]…”), while retaining the interleaved mask-text structure and keeping the number of mask tokens identical to the standard setting.
  2. 2.
[…]
  3. 3.
    Ablation on Benign Separators (w/o Separators): We removed all benign separators from the generated interleaved mask-text prompts, retaining only the vanilla harmful query followed by a block of mask tokens (i.e., Query + [MASK]…).
Ablations 2 and 3 essentially disrupt the critical interleaved mask-text prompt pattern.
Table 11: Ablation study results on HarmBench using LLaDA-8B-Instruct as the victim model. The high performance of w/o Refinement LLM confirms that the interleaved structure itself is the primary driver of the attack, while the poor performance of w/o Masking and w/o Separators highlights the necessity of both components.
Metric w/o Refinement LLM w/o Masking w/o Separators DiJA (Full)
ASR-k 94.5% 47.5% 53.5% 96.3%
ASR-e 54.8% 14.0% 16.8% 55.5%
Harmfulness Score (HS) 4.0 2.9 3.1 4.1
The experimental results, summarized in Table 11, reveal that the interleaved mask-text prompt structure is the key determinant of the attack’s success. Notably, even when the auxiliary LLM is removed (w/o Refinement LLM), the attack remains highly effective (ASR-e 54.8%), performing comparably to the full DiJA method (55.5%). This finding is critical as it rules out the hypothesis that our method relies on sophisticated prompt rewriting; rather, it confirms that the vulnerability is triggered by the specific structural constraints imposed on the dLLM. Conversely, removing either the masks (w/o Masking) or the separators (w/o Separators) leads to a drastic drop in performance, demonstrating that neither component works in isolation. This further highlights the fundamental difference between our method and prior rewriting-based jailbreak attacks: our approach is grounded in the intrinsic properties of dLLMs. By uncovering a new class of vulnerabilities specific to these models, we design an attack framework that is fundamentally aligned with their underlying mechanisms.
B.6 Impact of Mask Span Constraints
In real-world deployments, user interfaces might impose constraints on mask usage, such as limiting the number of allowed mask spans (e.g., only allowing a single “fill-in-the-blank” slot) or restricting the total length of masks. To understand the robustness of DiJA under such constraints, we conducted an additional ablation study. Specifically, we fixed the total budget of mask tokens to 50 and varied the number of separate mask spans allowed in the prompt from 1 to 10.
Table 12: Impact of the number of mask spans on attack success rate (LLaDA-8B-Instruct on HarmBench), with a fixed total mask budget of 50 tokens.
Number of Masked Spans 1 2 3 5 10
ASR-k 62.3% 85.1% 96.3% 94.8% 87.5%
ASR-e 31.5% 48.5% 55.5% 53.0% 43.5%
The results in Table 12 offer two key insights:
  1. 1.
    Necessity of Interleaving: Restricting the input to a single mask span ($N=1$) significantly degrades attack performance (ASR-e drops to 31.5%). This confirms that the interleaved mask-text structure, where text separators act as anchors between masks, is crucial for guiding the model’s generation and bypassing refusal mechanisms.
  2. 2.
    Trade-off with Span Length: While increasing the number of spans initially improves performance, excessive fragmentation (e.g., $N=10$) leads to a performance drop. This is because, with a fixed total budget, increasing the span count reduces the length of each individual span (e.g., 5 tokens per span), which may become too short to convey meaningful semantic content or instructions.
In summary, while strict UI constraints can serve as a partial mitigation, they do not fully eliminate the risk. DiJA remains potent as long as the interface allows for a moderate degree of interleaving.
B.7 Relation to Prefilling Attacks and Threat Model Assumptions
DiJA is conceptually related to prefilling attacks in autoregressive LLMs (Vega et al., 2023; Andriushchenko et al., 2024) in that both exploit the model’s tendency to preserve coherence and rely on the assumption that the user can exert control over the assistant-side output. However, DiJA generalizes these ideas to diffusion large language models (dLLMs), introducing unique capabilities that distinguish it from prior exploits on autoregressive models. Unlike prefilling attacks in autoregressive models, which are constrained to injecting adversarial prefixes due to the left-to-right nature of autoregressive decoding, DiJA leverages the non-autoregressive property (i.e., parallel decoding and bidirectional context modeling) of dLLMs to manipulate arbitrary masked spans within the assistant response. By performing iterative span-level rewriting rather than relying solely on prefix injection, DiJA achieves more flexible and fine-grained control over generated outputs. It is important to emphasize that, beyond merely proposing an attack technique targeting dLLMs, our work primarily aims to expose emergent security vulnerabilities that arise from the unique characteristics of diffusion-based language models. By doing so, we hope to draw the community’s attention to these novel risks and provide valuable insights for the development and safety alignment of future dLLM training paradigms.
Regarding the threat model, DiJA assumes the ability to perform remasking or editing on the assistant-side response. This condition is trivially satisfied in open-source models where users have full access to the model weights. For closed-source models, the attack requires the provider to expose an editing or infilling API; without such features, DiJA can be mitigated by disabling arbitrary mask placement. We draw a parallel to the autoregressive setting: while OpenAI’s API prevents prefilling attacks by restricting user control over the assistant output, Anthropic’s Claude API [^10] https://anthropic.mintlify.app/en/docs/build-with-claude/prompt-engineering/prefill-claudes-response explicitly supports prefilling, thereby satisfying the conditions for such attacks (Andriushchenko et al., 2024). Similarly, while commercial dLLM APIs are currently nascent, remasking and re-generating is a core feature of dLLMs and has a valuable application prospect (as shown in Figure 1). Thus, future deployments exposing this native capability will naturally fall under the threat model of DiJA.
Appendix C More Implementation Details on DiJA
C.1 Victim Models